OpenStudy (macaylam):

Does a two-digit number exist such that the digits sum to 9 and when the digits are reversed the resulting number is 9 greater than the original number? Identify the system of equations that models the given scenario. t + u = 9 10t + u = 10u + t – 9 t + u = 9 10t + u = 10u + t t + u = 9 tu = ut + 9

OpenStudy (mathstudent55):

t = tens digit u = units digit

OpenStudy (mathstudent55):

The original number is 10t + u Ok so far?

OpenStudy (macaylam):

yeah

OpenStudy (mathstudent55):

The original number is actually 10t + u. The sum of its digits is simply t + u Ok?

OpenStudy (mathstudent55):

The sum of the digits of the original number is 9, so we have t + u = 9 That is one equation. Now we need the second equation.

OpenStudy (mathstudent55):

The original number is 10t + u The new has the digits reversed, so the new number is 10u + t

OpenStudy (mathstudent55):

The new number is 9 greater than the original number. That means 10u + t is 9 greater than 10t + u or 10u + t = 10t + u + 9

OpenStudy (mathstudent55):

Subtract 9 from both sides of the last equation: 10u + t - 9 = 10t + u Switch sides: 10t + u = 10u + t - 9 This is our second equation. The system of equations is: t + u = 9 10t + u = 10u + t - 9 Which option is that?

OpenStudy (macaylam):

A

OpenStudy (mathstudent55):

Correct.

OpenStudy (macaylam):

thanks for helping (:

OpenStudy (mathstudent55):

You're welcome.